OpenStudy (anonymous):

The area of a square garden is 128 m². How long is the diagonal? A. 8 m B. 4 times the square root of 2 C. eight times the square root of 6 D. 16 m My guess is D

OpenStudy (campbell_st):

well take the square root of the area to find the side length. then use pythagoras to find the diagonal length

OpenStudy (anonymous):

\[S ^{2}=128\] Solve for S and the use the pythagorean to find the diagonal

OpenStudy (anonymous):

\[\sqrt{2S ^{2}}=D\]

OpenStudy (anonymous):

Not to sound moronic but whats pythagoras?

OpenStudy (anonymous):

Like is it a formula?

OpenStudy (anonymous):

pythagoras is a mathematician who developed the pythagorean theorem to solve for the hypotenuse of a right triangle. the formula you will use is of the form \[a ^{2}+b ^{2}=c ^{2}\] where a and b are the sides of the legs. I have reduced the equation further for your particular case with the equation above

OpenStudy (campbell_st):

ok... to start you need to know  using thr diagram and applying pythagoras' theorem to find the diagonal \[d^2 = l^2 + l^2\] which becomes \[d^2 = 128 + 128\] can you simplify this?